# Lifelong Knowledge Editing for LLMs with Retrieval-Augmented Continuous Prompt Learning

### v2

@@ Abstract
Model editing aims to correct outdated or erroneous knowledge in large language models (LLMs) without the need for costly retraining. Lifelong model editing is the most challenging task that caters to the continuous editing requirements of LLMs. Prior works primarily focus on single or batch editing; nevertheless, these methods fall short in lifelong editing scenarios due to catastrophic knowledge forgetting and the degradation of model performance. Although retrieval-based methods alleviate these issues, they are impeded by slow and cumbersome processes of integrating the retrieved knowledge into the model. In this work, we introduce RECIPE, a RetriEval-augmented ContInuous Prompt lEarning method, to boost editing efficacy and inference efficiency in lifelong learning. RECIPE first converts knowledge statements into short and informative continuous prompts, prefixed to the LLM’s input query embedding, to efficiently refine the response grounded on the knowledge. It further integrates the Knowledge Sentinel (KS) that acts as an intermediary to calculate a dynamic threshold, determining whether the retrieval repository contains relevant knowledge. Our retriever and prompt encoder are jointly trained to achieve editing properties, i.e., reliability, generality, and locality. In our experiments, RECIPE is assessed extensively across multiple LLMs and editing datasets, where it achieves superior editing performance. RECIPE also demonstrates its capability to maintain the overall performance of LLMs alongside showcasing fast editing and inference speed.  [^1] Source codes will be released upon paper acceptance.
@@ 1 Introduction
Previous efforts in model editing have primarily focused on single and batch edits. Notable examples include ROME Meng et al. (2022), MEND Mitchell et al. (2022), and MEMIT Meng et al. (2023), which achieve edits by applying offsets to part of the model’s parameters. However, in the real world, LLMs frequently require continuous knowledge updates to stay abreast of emerging knowledge. Thus, the concept of lifelong editing has been introduced Hartvigsen et al. (2022). As shown in the upper part of Figure 1, with continuous editing, the accumulating offsets on parameters can result in model performance degradation or even failure Hartvigsen et al. (2022); Huang et al. (2023); Han et al. (2023); Hu et al. (2024).
Some techniques Dong et al. (2022); Huang et al. (2023) address the challenges by integrating extra model parameters. Nevertheless, as shown in the middle of Figure 1, the increase in additional parameters leads to diminished model performance and reduced inference efficiency. Retrieval-based methods Han et al. (2023); Jiang et al. (2024); Yu et al. (2024) separate knowledge from the model, thereby alleviating knowledge forgetting and performance degradation. However, the intricate post-retrieval knowledge adoption inevitably reduces the inference efficiency of LLMs, such as appending a lengthy knowledge updating instruction before the input query Jiang et al. (2024).
@@
Knowledgeable Continuous Prompt Learning: In RECIPE, each editing knowledge statement (expressed by texts) is transformed into a knowledgeable continuous prompt. The prompt is then prefixed before the embedding of the input query to modify the response of the LLM. This approach is grounded in prior research, as exemplified by P-tuning Li and Liang (2021); Liu et al. (2022), which demonstrated that continuous prompts enable LLMs to perform downstream tasks more effectively. Here, we conceptualize each knowledge edit as a distinct mini-task. To ensure editing efficacy, our prompt encoder is trained to align with three key editing properties including reliability, generality, and locality Yao et al. (2023).
Dynamic Prompt Retrieval with Knowledge Sentinel: Initially, we map the knowledge statements and queries into the same representational space to compute retrieval similarity. Manually setting a fixed similarity threshold is a common practice to determine whether the repository contains knowledge related to an input query Han et al. (2023). However, this approach does not account for the fact that different queries often require distinct thresholds due to semantic variations. Therefore, we introduce the Knowledge Sentinel (KS), a trainable embedding representation, as an intermediary to dynamically compute the threshold for each query. Employing a specifically designed contrastive learning mechanism, the KS module is jointly trained with the prompt encoder to align retrieval with model editing.
In the experiments, we conduct tests with 1, 10, 100, 1,000, and 10,000 edits to compare the performance of our model against prominent editing methods on ZSRE Mitchell et al. (2022), CounterFact Meng et al. (2022), and RIPE Cohen et al. (2023) using LLaMA-2 (7B), GPT-J (6B) and GPT2-XL (1.5B) backbones. Results demonstrate that RECIPE achieves not only optimal editing performance and robustness against degradation of LLM general results but also a significant advantage in both editing and inference speed.
@@ 2.1 Model Editing
Methods modifying model parameters can be further divided into Locate-then-Edit (L&E) and meta-learning-based methods. For L&E, ROME Meng et al. (2022) identifies the LLMs’ edit-sensitive layers through causal tracing and proposes rank-one model editing to modify parameters. MEMIT Meng et al. (2023) and WILKE Hu et al. (2024) respectively use multi-layer allocation and dynamic localization to alleviate the single matrix update burden of ROME. In meta-learning-based methods, KnowledgeEditor Cao et al. (2021) and MEND Mitchell et al. (2022) respectively transform editing knowledge and the gradient decomposition of LLM to the offsets of the weights to be edited. MALMEN Tan et al. (2023) enhances MEND’s approach by using normal equations to merge parameters for multiple edits. Although these methods show success in single or batch editing scenarios, in a lifelong editing situation, as the number of edits increases, the accumulating mismatches of parameter offsets can lead to model degradation or failure Hu et al. (2024).
@@
Retrieval-based editors effectively circumvent the issue of accumulated parameter offsets and the potentially unbounded addition of neurons. GRACE Hartvigsen et al. (2022) maintains an adapter that maps a query to a potential representation corresponding to the knowledge retrieved by calculating the distance between representations of the query and the knowledge. RASE Han et al. (2023) develops an editing retrieval model to boost the efficacy of model editing approaches during sequential edits. Complementing GRACE, MELO Yu et al. (2024) introduces a batch editing version using LoRA Hu et al. (2022). LTE Jiang et al. (2024) fine-tunes the LLM to respond to knowledge when prefixed with editing information and retrieves relevant content using the off-the-shelf backbone Reimers and Gurevych (2019). While retrieval-based methods are advantageous for lifelong learning, they may still contend with speed issues and practical complexities involved in editing and applying knowledge after retrieval.
@@ 4.3 Model Inference with Editing On-the-fly
# Editing Editor ZSRE CF RIPE
Rel. Gen. Loc. Avg. Rel. Gen. Loc. Avg. Rel. Gen. Loc. Avg.
1 FT 47.86 42.57 93.89 61.44(±1.00) 41.37 26.04 52.25 39.89(±0.74) 41.54 33.89 53.27 42.90(±0.33)
TP 86.35 83.98 86.34 85.56(±0.53) 91.41 68.61 38.94 66.32(±1.18) 76.98 55.10 51.29 61.13(±0.48)
GRACE 99.20 33.23 99.82 77.42(±0.78) 98.65 11.42 98.73 69.60(±0.66) 98.13 28.45 99.75 75.44(±0.65)
10 FT 44.08 43.98 70.03 52.70(±0.30) 18.09 15.49 21.53 18.37(±1.39) 22.74 18.51 18.84 20.03(±0.53)
TP 57.33 52.37 36.68 48.79(±1.47) 85.92 58.64 21.56 55.37(±0.43) 63.38 41.20 30.45 45.01(±0.80)
GRACE 52.10 36.64 98.80 62.51(±0.78) 60.61 11.89 96.52 56.34(±0.78) 49.14 30.68 98.30 59.37(±0.96)
100 FT 35.95 33.35 25.30 31.54(±1.06) 3.63 0.19 0.01 1.27(±0.32) 9.36 5.19 6.19 6.91(±1.00)
TP 46.05 41.20 9.67 32.30(±0.91) 70.01 40.76 4.51 38.42(±0.54) 44.73 28.94 11.60 28.42(±0.91)
GRACE 47.62 34.99 98.00 60.21(±0.41) 55.00 12.85 93.49 53.78(±0.54) 41.03 31.02 95.15 55.74(±0.44)
1000 FT 14.66 12.61 2.69 9.99(±1.00) 6.94 0.68 3.48 3.70(±0.09) 7.91 2.13 1.82 3.95(±0.40)
TP 44.72 41.38 4.38 30.16(±1.04) 64.70 32.50 11.63 36.28(±0.72) 42.24 26.80 9.87 26.30(±1.01)
GRACE 42.04 33.42 96.73 57.40(±0.68) 52.75 12.86 91.02 52.21(±0.85) 38.03 30.10 91.24 53.12(±0.61)
10000 FT 6.57 5.29 0.44 4.10(±0.24) 4.86 0.76 2.19 2.60(±1.40) - - - -
MEND 0.01 0.02 0.01 0.01(±0.01) 0.03 0.01 0.00 0.01(±0.00) - - - -
ROME 1.21 0.38 0.15 0.58(±0.19) 0.16 0.07 0.22 0.15(±0.07) - - - -
MEMIT 0.03 0.01 0.02 0.02(±0.00) 0.03 0.02 0.03 0.02(±0.01) - - - -
TP 37.53 33.55 3.94 25.01(±0.56) 58.26 29.25 11.42 32.98(±0.67) - - - -
GRACE 38.50 31.52 93.15 54.39(±0.43) 48.52 11.75 85.38 48.55(±1.78) - - - -
WILKE 7.39 5.11 14.05 8.85(±1.15) 5.17 3.70 23.87 10.91(±1.17) - - - -
Table 1: The overall results using LLAMA-2 (7B) in lifelong editing scenario. Due to the space limitation, our editing results of GPT-J and GPT-XL are shown in Appendix D. “# Editing” denotes the number of edits. “Rel.”, “Gen.” and “Loc.” are the abbreviations of reliability, generality, and locality, respectively. Given that the RIPE dataset comprises 4,388 samples, achieving results for 10,000 edits is not feasible. The t-tests demonstrate the improvements of our work are statistically significant with $p$ < 0.05 level.
@@ 4.4 Model Training
where $p_{k_{i}}$ is the continuous prompt transformed through Eq. 6 and Eq. 7 using knowledge $k_{i}$ that is the concatenation of $q_{e_{i}}$ and $a_{e_{i}}$. The $\mathrm{KL}$ denotes the Kullback-Leibler divergence. The batch-wise loss function for model editing is derived as follows:
@@
Table 3: Average time (s) taken for a single edit and model inference after 10,000 edits. MP, AP, and RB indicate Modifying Parameters, Adding Parameters, and Retrieval-Based methods, respectively.
@@ 5 Experiments
In this section, we present the experimental results of RECIPE and compare it against strong baselines over various public datasets  [^2] The detailed description of datasets, baselines, and model settings are presented in Appendix B and Appendix C..
@@ 5.1 The Performance of RECIPE
We evaluate RECIPE using various backbones including LLAMA-2 (7B), GPT-J (6B), and GPT-XL (1.5B) shown in Table 1 and Appendix D.
Editing Performance: Table 1 presents the overall performance across various numbers of edits to simulate a lifelong editing scenario. From the single-edit perspective, our method exhibits optimal performance in most testing scenarios. In the lifelong editing scenarios, we have the following observations: (1) Methods that modify the parameters of LLMs, e.g., MEND Mitchell et al. (2022), ROME Meng et al. (2022), and MEMIT Meng et al. (2023)), show outstanding editing performance in a single edit. Yet, they exhibit a significant decline in editing performance as the number of edits increases. This trend aligns with the toxic accumulation issue highlighted by Hu et al. (2024). (2) Methods introducing additional parameters, such as T-Patcher Huang et al. (2023), maintain a degree of reliability and generality in the lifelong editing process. However, the cumulative addition of extra parameters compromises the original inference process, evidenced by the pronounced deterioration in locality observed in ZSRE Mitchell et al. (2022). (3) Retrieval-based approaches, including GRACE Hartvigsen et al. (2022), R-ROME Han et al. (2023), and LTE Jiang et al. (2024), demonstrate robustness against the increasing number of edits. Our method achieves the best results, affirming the strengths of retrieval as well as validating the efficacy of our strategy.
Damage to the General Performance of LLMs: While the three editing metrics effectively demonstrate the editing performance, we further investigate to which extent these editors influence the model’s general capabilities. Table 2 shows the results of LLaMA-2 after 1,000 edits. It is observed that non-retrieval-based methods lead to a significant reduction in general capabilities. This can be attributed to the accumulation of pattern mismatches caused by external interventions of editing. Among retrieval-based methods, LTE Jiang et al. (2024) also exhibits performance degradation. In contrast, our RECIPE does not involve direct intervention on LLM parameters but instead relies on concatenating a short prompt to guide the LLM’s adherence to knowledge. It demonstrates the best preservation of general performance, suggesting that it inflicts minimal harm on the model.

### v3

@@ Abstract
Model editing aims to correct outdated or erroneous knowledge in large language models (LLMs) without the need for costly retraining. Lifelong model editing is the most challenging task that caters to the continuous editing requirements of LLMs. Prior works primarily focus on single or batch editing; nevertheless, these methods fall short in lifelong editing scenarios due to catastrophic knowledge forgetting and the degradation of model performance. Although retrieval-based methods alleviate these issues, they are impeded by slow and cumbersome processes of integrating the retrieved knowledge into the model. In this work, we introduce RECIPE, a RetriEval-augmented ContInuous Prompt lEarning method, to boost editing efficacy and inference efficiency in lifelong learning. RECIPE first converts knowledge statements into short and informative continuous prompts, prefixed to the LLM’s input query embedding, to efficiently refine the response grounded on the knowledge. It further integrates the Knowledge Sentinel (KS) that acts as an intermediary to calculate a dynamic threshold, determining whether the retrieval repository contains relevant knowledge. Our retriever and prompt encoder are jointly trained to achieve editing properties, i.e., reliability, generality, and locality. In our experiments, RECIPE is assessed extensively across multiple LLMs and editing datasets, where it achieves superior editing performance. RECIPE also demonstrates its capability to maintain the overall performance of LLMs alongside showcasing fast editing and inference speed.  [^1] Source codes is available at https://github.com/qizhou000/RECIPE.
Lifelong Knowledge Editing for LLMs with Retrieval-Augmented Continuous Prompt Learning
Qizhou Chen1,2thanks:   Q. Chen and T. Zhang contributed equally to this work., Taolin Zhang211footnotemark: 1, Xiaofeng He1,3, Dongyang Li1,2, Chengyu Wang4thanks:   Corresponding author., Longtao Huang2, Hui Xue2 1 East China Normal University, Shanghai, China 2 Alibaba Group, Hangzhou, China 3NPPA Key Laboratory of Publishing Integration Development, ECNUP, Shanghai, China 4 Alibaba Cloud Computing, Hangzhou, China chen_qizhou@outlook.com, zhangtl0519@gmail.com, chengyu.wcy@alibaba-inc.com
@@ 1 Introduction
Previous efforts in model editing have primarily focused on single and batch edits. Notable examples include ROME Meng et al. (2022), MEND Mitchell et al. (2022), and MEMIT Meng et al. (2023), which achieve edits by applying offsets to part of the model’s parameters. However, in the real world, LLMs frequently require continuous knowledge updates to stay abreast of emerging knowledge. Thus, the concept of lifelong editing has been introduced Hartvigsen et al. (2022). As shown in the upper part of Figure 1, with continuous editing, the accumulating offsets on parameters can result in model performance degradation or even failure Hartvigsen et al. (2022); Huang et al. (2023); Han et al. (2023); Hu et al. (2024). Some techniques Dong et al. (2022); Huang et al. (2023) address the challenges by integrating extra model parameters. Nevertheless, as shown in the middle of Figure 1, the increase in additional parameters leads to diminished model performance and reduced inference efficiency.
Retrieval-based methods separate knowledge from the model. As shown in the lower part of Figure 1, they temporarily integrate retrieved knowledge into the model during each inference, mitigating knowledge forgetting and performance degradation. However, these methods face certain limitations: retrieval at each token prediction Hartvigsen et al. (2022), caching large update matrices Han et al. (2023); Yu et al. (2024), and overly long editing prefixes Jiang et al. (2024). These shortcomings affect the model’s inference efficiency.
@@
Knowledgeable Continuous Prompt Learning: In retrieval-based methods, LTE Jiang et al. (2024) effectively avoids the shortcomings of earlier approaches. However, their overly long editing prefixes can reduce model inference speed, and full-parameter fine-tuning also increases the risk of overfitting. Therefore, we aim to explore the shortest possible prefixes to enable the LLM to follow editing instructions. In RECIPE, we achieve this by constructing a continuous short prompt encoder that transforms each piece of editing knowledge (expressed as text) into a knowledgeable continuous prompt. This approach is grounded in prior research, as exemplified by P-tuning Li and Liang (2021); Liu et al. (2022), which demonstrated that continuous prompts enable LLMs to perform downstream tasks more effectively. Here, we conceptualize each knowledge edit as a distinct mini-task. To ensure editing efficacy, our prompt encoder is trained to align with three key editing properties including reliability, generality, and locality Yao et al. (2023).
Dynamic Prompt Retrieval with Knowledge Sentinel: To avoid unnecessary costs from introducing additional retrievers Han et al. (2023); Jiang et al. (2024), we map knowledge statements and queries into the same representational space through the prompt encoder to compute retrieval similarity. To determine whether the retrieval repository contains knowledge related to an input query, manually setting a fixed similarity threshold is a common practice Han et al. (2023). However, this method does not take into account that different queries often require distinct thresholds due to semantic variations. Therefore, we introduce the Knowledge Sentinel (KS), a trainable embedding representation, as an intermediary to dynamically compute the threshold for each query. Employing a specifically designed contrastive learning mechanism, the KS module is jointly trained with the prompt encoder to align retrieval with model editing.
The contributions of our paper are summarized as follows:
  • We propose a novel RetriEval-augmented Continuous Prompt Learning framework, RECIPE, to facilitate lifelong model editing. As far as we know, we are the first to adapt prompt learning for model editing, and explore the shortest possible editing prefixes.
  • Within RECIPE, the prompt learning encoder effectively shortens the editing prefixes, enhancing the inference efficiency of the post-edit LLM. The KS improves editing retrieval, thus enhancing overall editing performance.
  • We conduct comparative experiments of lifelong editing across multiple backbones and editing datasets. The results demonstrate the superiority of RECIPE.
@@ 2.1 Model Editing
Methods modifying model parameters can be further divided into Locate-then-Edit (L&E) and meta-learning-based methods. For L&E, ROME Meng et al. (2022) identifies the LLMs’ edit-sensitive layers through causal tracing and proposes rank-one model editing to modify parameters. MEMIT Meng et al. (2023) and WILKE Hu et al. (2024) respectively use multi-layer allocation and dynamic localization to alleviate the single matrix update burden of ROME. In meta-learning-based methods, KnowledgeEditor Cao et al. (2021) and MEND Mitchell et al. (2022) respectively transform editing knowledge and the gradient decomposition of LLM to the offsets of the weights to be edited. MALMEN Tan et al. (2023) and DAFNet Zhang et al. (2024) enhance MEND’s parameter fusion in multiple editing by utilizing normal equations and constructing interactions between edits, respectively. Although these methods show success in single or batch editing scenarios, in a lifelong editing situation, as the number of edits increases, the accumulating mismatches of parameter offsets can lead to model degradation or failure Hu et al. (2024).
@@
Retrieval-based editors effectively circumvent the issue of accumulated parameter offsets and the potentially unbounded addition of neurons. GRACE Hartvigsen et al. (2022) performs editing retrieval for each token prediction by calculating the linear distance between representations, which impacts the efficiency of long sequence predictions. RASE Han et al. (2023) develops an editing retrieval model to improve sequential editing. MELO Yu et al. (2024) introduces a batch editing version of GRACE using LoRA Hu et al. (2022). Both methods’ instant editing schemes require updating the corresponding FFN matrices separately for each sample in the input batch, which increases memory load as the batch size grows. LTE Jiang et al. (2024) fine-tunes the LLM to respond to knowledge when prefixed with editing information and retrieves relevant content using the off-the-shelf backbone Reimers and Gurevych (2019). However, lengthy editing prefixes still impact inference efficiency. Additionally, fine-tuning the model itself is likely to lead to overfitting, thereby affecting its original performance. To address the above issues, I propose RECIPE to explore more efficient retrieval and instant editing, as well as smaller interventions to the model to avoid overfitting on the editing dataset.
@@ 4.4 Model Training
where $p_{k_{i}}$ is the continuous prompt transformed through Eq. 6 and Eq. 7 using knowledge $k_{i}$ that is the concatenation of $q_{e_{i}}$ and $a_{e_{i}}$. The $\mathrm{KL}$ denotes the Kullback-Leibler divergence, which is chosen to better fit the LLM’s original prediction distribution on the locality data. The batch-wise loss function for model editing is derived as follows:
@@
# Editing Type Editor ZSRE CF RIPE
Rel. Gen. Loc. Avg. Rel. Gen. Loc. Avg. Rel. Gen. Loc. Avg.
1 MP FT 47.86 42.57 93.89 61.44(±1.00) 41.37 26.04 52.25 39.89(±0.74) 41.54 33.89 53.27 42.90(±0.33)
AP TP 86.35 83.98 86.34 85.56(±0.53) 91.41 68.61 38.94 66.32(±1.18) 76.98 55.10 51.29 61.13(±0.48)
RB GRACE 99.20 33.23 99.82 77.42(±0.78) 98.65 11.42 98.73 69.60(±0.66) 98.13 28.45 99.75 75.44(±0.65)
10 MP FT 44.08 43.98 70.03 52.70(±0.30) 18.09 15.49 21.53 18.37(±1.39) 22.74 18.51 18.84 20.03(±0.53)
AP TP 57.33 52.37 36.68 48.79(±1.47) 85.92 58.64 21.56 55.37(±0.43) 63.38 41.20 30.45 45.01(±0.80)
RB GRACE 52.10 36.64 98.80 62.51(±0.78) 60.61 11.89 96.52 56.34(±0.78) 49.14 30.68 98.30 59.37(±0.96)
100 MP FT 35.95 33.35 25.30 31.54(±1.06) 3.63 0.19 0.01 1.27(±0.32) 9.36 5.19 6.19 6.91(±1.00)
AP TP 46.05 41.20 9.67 32.30(±0.91) 70.01 40.76 4.51 38.42(±0.54) 44.73 28.94 11.60 28.42(±0.91)
RB GRACE 47.62 34.99 98.00 60.21(±0.41) 55.00 12.85 93.49 53.78(±0.54) 41.03 31.02 95.15 55.74(±0.44)
1000 MP FT 14.66 12.61 2.69 9.99(±1.00) 6.94 0.68 3.48 3.70(±0.09) 7.91 2.13 1.82 3.95(±0.40)
AP TP 44.72 41.38 4.38 30.16(±1.04) 64.70 32.50 11.63 36.28(±0.72) 42.24 26.80 9.87 26.30(±1.01)
RB GRACE 42.04 33.42 96.73 57.40(±0.68) 52.75 12.86 91.02 52.21(±0.85) 38.03 30.10 91.24 53.12(±0.61)
10000 MP FT 6.57 5.29 0.44 4.10(±0.24) 4.86 0.76 2.19 2.60(±1.40) - - - -
MEND 0.01 0.02 0.01 0.01(±0.01) 0.03 0.01 0.00 0.01(±0.00) - - - -
ROME 1.21 0.38 0.15 0.58(±0.19) 0.16 0.07 0.22 0.15(±0.07) - - - -
MEMIT 0.03 0.01 0.02 0.02(±0.00) 0.03 0.02 0.03 0.02(±0.01) - - - -
WILKE 7.39 5.11 14.05 8.85(±1.15) 5.17 3.70 23.87 10.91(±1.17) - - - -
AP TP 37.53 33.55 3.94 25.01(±0.56) 58.26 29.25 11.42 32.98(±0.67) - - - -
RB GRACE 38.50 31.52 93.15 54.39(±0.43) 48.52 11.75 85.38 48.55(±1.78) - - - -
Table 1: The overall results using LLAMA-2 (7B) in lifelong editing scenario. Editing results of GPT-J and GPT-XL are shown in Appendix D. “# Editing” denotes the number of edits. “Rel.”, “Gen.” and “Loc.” are the abbreviations of reliability, generality, and locality, respectively. Given that the RIPE dataset comprises 4,388 samples, achieving results for 10,000 edits is not feasible. MP, AP, and RB indicate Modifying Parameters, Adding Parameters, and Retrieval-Based methods, respectively. The t-tests demonstrate the improvements of our work are statistically significant with $p$ < 0.05 level.
@@
Table 3: Average time (s) taken for a single edit and model inference after 10,000 edits.
@@ 5 Experiments
In this section, we present the experimental results on LLMs including LLAMA-2 (7B), GPT-J (6B), and GPT-XL (1.5B). We compare RECIPE against strong baselines including Fine-Tune (FT) Yao et al. (2023), MEND Mitchell et al. (2022), ROME Meng et al. (2022), MEMIT Meng et al. (2023), MALMEN Tan et al. (2023), WILKE Hu et al. (2024), T-Patcher (TP) Huang et al. (2023), GRACE Hartvigsen et al. (2022), R-ROME Han et al. (2023), and LTE Jiang et al. (2024). For the evaluation of lifelong editing, we use datasets including ZSRE Mitchell et al. (2022), CF Meng et al. (2022), and RIPE Cohen et al. (2023). For the evaluation of general performance after editing, we apply datasets including CSQA Talmor et al. (2019), ANLI Nie et al. (2020), MMLU Hendrycks et al. (2021), and SQuAD-2 Rajpurkar et al. (2018). For more detailed description of datasets, baselines, and model settings, please refer to Appendix B and Appendix C.
@@ 5.1 The Performance of RECIPE
Editing Performance: Table 1 and Appendix D present the overall editing performance across various numbers of edits to simulate a lifelong editing scenario. From the single-edit perspective, our method exhibits optimal performance in most testing scenarios. In the lifelong editing scenarios, we have the following observations: (1) Methods that modify the parameters of LLMs show outstanding editing performance in a single edit. Yet, they exhibit a significant decline in editing performance as the number of edits increases. This trend aligns with the toxic accumulation issue highlighted by Hu et al. (2024). (2) Method introducing additional parameters maintains a degree of reliability and generality in the lifelong editing process. However, the cumulative addition of extra parameters compromises the original inference process, evidenced by the pronounced deterioration in locality observed in ZSRE. (3) Retrieval-based approaches demonstrate robustness against the increasing number of edits. Among them, our method achieves the best results, affirming the strengths of retrieval as well as validating the efficacy of our strategy.
From the perspective of metrics, for reliability and generality, the co-occurrence of correct retrieval and the model’s faithful adherence to editing instructions is a prerequisite for editing success. Therefore, compared to other retrieval-based methods, we consider the additional introduction of KS in RECIPE, which enhances retrieval performance and thereby improves editing efficacy, to be a crucial factor for RECIPE’s superior performance in terms of reliability and generality metrics, as demonstrated in Table 4. Regarding locality, aside from the ability of the editing retrieval to effectively exclude irrelevant knowledge, the locality component of the editing loss for RECIPE, along with the shorter continuous prompt editing prefixes, also minimizes the impact on the model’s editing-unrelated responses.
Damage to the General Performance of LLMs: While the three editing metrics effectively demonstrate the editing performance, we further investigate to which extent these editors influence the model’s general capabilities. Table 2 shows the results of LLaMA-2 after 1,000 edits. It is observed that non-retrieval-based methods lead to a significant reduction in general capabilities. This can be attributed to the accumulation of pattern mismatches caused by external interventions of editing. Among retrieval-based methods, LTE also exhibits performance degradation. In contrast, our RECIPE does not involve direct intervention on LLM parameters but instead relies on concatenating a short prompt to guide the LLM’s adherence to knowledge. It demonstrates the best preservation of general performance, suggesting that it inflicts minimal harm on the model.